# Combinatorial Creativity: A New Frontier in Generalization Abilities

### v1

@@ Abstract
Artificial intelligence (AI) systems, and large language models (LLMs) in particular, are increasingly employed for creative tasks like scientific idea generation, constituting a form of generalization from training data unaddressed by existing conceptual frameworks. Though in many ways similar to forms of compositional generalization (CG), combinatorial creativity (CC) is an open-ended ability. Instead of evaluating for accuracy or correctness against fixed targets, which would contradict the open-ended nature of CC, we propose a theoretical framework and algorithmic task for evaluating outputs by their degrees of novelty and utility. From here, we make several important empirical contributions: (1) We obtain the first insights into the scaling behavior of creativity for LLMs. (2) We discover that, for fixed compute budgets, there exist optimal model depths and widths for creative ability. (3) We find that the ideation-execution gap, whereby LLMs excel at generating novel scientific ideas but struggle to ensure their practical feasibility, may be explained by a more fundamental novelty-utility tradeoff characteristic of creativity algorithms in general. Importantly, this tradeoff remains persistent even at scale, casting doubt on the long-term creative potential of LLMs in their current form. Together, our conceptual framework and empirical findings provide a foundation for understanding and improving creativity in modern AI models, marking a new frontier in generalization abilities.
@@ 1 Introduction
Modern AI systems now engage in scientifically creative tasks once conceptualized by Einstein (Gu and Krenn, 2024, Si et al., 2024, Sanyal et al., 2025), yet they lack strong mathematical and conceptual foundations for the abilities underlying these tasks. As a result, many problems have surfaced. LLM-generated ideas for scientific discovery often suffer from practical infeasibility, make unrealistic assumptions, and omit proper baselines, leading to what has been termed the ideation-execution gap (Si et al., 2025). Without a foundational understanding of creativity, our ability to diagnose and improve the outcomes of LLMs for such tasks remains severely limited.
To address these limitations, we introduce a formal framework and an open-ended, algorithmic task for evaluating combinatorial creativity. Our framework models creativity within a conceptual space, represented as a large graph where models must find novel paths between concepts while adhering to logical constraints. Using this testbed, we conduct a large-scale empirical study of decoder-only Transformers, systematically varying their size, depth, and width to yield several key insights.
First, we obtain the first insights into the scaling behavior for creativity, demonstrating a predictable improvement in performance with increased model size and training compute. Second, we uncover a crucial architectural insight: for a fixed computational budget, wider, shallower architectures consistently outperform their deeper, narrower counterparts. However, this preference for width is not limitless; we identify an optimal aspect ratio (width-to-depth) that maximizes creative performance, beyond which returns diminish. Third, we perform a detailed error analysis, which reveals that as task complexity increases, models are more likely to fail by violating utility constraints rather than by producing non-novel outputs. This is explained by our fourth and most critical finding: a fundamental novelty-utility tradeoff, which persists even as model scale increases. This suggests the ideation-execution gap may be an intrinsic challenge for current architectures; not merely a problem of scale. Together, our conceptual framework and empirical findings offer a launchpad for understanding and improving the creativity of modern AI models, as we enter a new frontier in generalization abilities.
@@ 2.2 Distinguishing Combinatorial Creativity from Classical Forms of Generalization
In Table 1, we compare generalization abilities along six key aspects. An ability is compositional (A1) if it involves recombination of atomic units into compound artifacts; open-ended [^1] Note that our notion of open-endedness is slightly different from the recent definition in Hughes et al. (2024) because we consider open-endedness from the product, not process, perspective (Rhodes, 1961) (A2) if there is no single correct answer for its evaluation, but instead multiple plausible answers; structurally novel (A3) if it generates artifacts whose form is distinct from structures trained on; and semantically novel (A4) if generated artifacts have new meanings. Lastly, an ability involves measuring degrees of novelty (A5) and degrees of utility (A6) if artifacts may be more or less novel or useful, respectively, depending on their semantic or structural properties.
@@ 3.2 Quantifying Degrees of Novelty
In order for an artifact to be judged creative, one condition is that it must be novel. Given an artifact $P$, there are two common ways to measure its novelty: (i) as some function of the distance $d$ between $P$ and a set of existing artifacts $d(f(P))$ (Maher, 2010), or, for combinatorial creativity especially, (ii) semantic graph distances induced by the combinatorial components [^2] Note that under certain conditions, semantic graph distances are asymptotically equivalent to statistical distances to existing artifact sets (Varshney et al., 2020) (Varshney et al., 2020, Gray et al., 2019). To keep the algorithmic task as controllable as possible, we adopt method (ii), quantifying novelty via the graph walk distance and the surprise of the labels used on the walk, which can be understood as a proxy for semantic distance (Gray et al., 2019).
@@ 3.3 Quantifying Degrees of Utility
The utility function consists of three main parts: the terms $\left(1+\alpha_{I}|\mathcal{I}|\right)$ and $\left(1+\alpha_{X}|\mathcal{X}|\right)$ scale the utility function in proportion to the number of inclusion and exclusion constraints, respectively, while the indicator term ensures that artifacts obey these constraints and start and end at the correct nodes. The number of total constraints also allows us to informally define “utility levels:” level $1$ consists of an unconstrained query ($I=\emptyset,X=\emptyset$), while levels $l>1$ each consist of l1l-1 constraints. In Section 5, we use these utility levels to demonstrate the novelty-utility tradeoff in creative LLM outputs, and we also explain how the constraints can serve as minimal abstractions of failure modes encountered when using LLMs for scientifically creative idea generation.
@@ 4 Experiments
From here, we are interested in how fundamental architectural choices influence the creativity of LLMs on the task defined in Section 3. For example, Nagarajan et al. (2025) recently found creative gains from changing the pre-training objective from next-token to multi-token prediction. In this study, we are especially curious whether creative limits present in smaller models also persist at scale, as this can provide insight into whether continued scaling of LLMs can yield even greater gains in creativity, or if architectural innovations are needed to support more open-ended behaviors (Hughes et al., 2024).
@@ 5 Results and Discussion
Figure 3: The impact of width and depth on creativity. These heatmaps visualize the combinatorial creativity of models across three distinct parameter budgets (1M, 10M, and 100M). For each budget, the vertical axis represents the amount of training compute in FLOPs. The color intensity corresponds to the model’s creativity score, while the horizontal axis represents (top) the number of layers $L$ or (bottom) the width to depth ratio $E/L$. The contours reveal a clear, non-monotonic trend: (top) creativity improves as layers are added up to a certain point, after which performance declines, and creativity improves as the width is increased up to a certain point, after which performance also declines (bottom). The optimal depth becomes more pronounced at larger scales, with the 100M models achieving peak creativity around 8 layers, while the optimal performance for width is at an $E/L$ ratio between 200 and 300.
@@
In Figure 3 (top), we visualize the impact of the number of layers $L$ on combinatorial creativity across all three model sizes. Our most significant finding is that for a fixed parameter count, there is an architectural “sweet spot,” an optimal number of layers that maximizes creativity, after which increasing depth further can be detrimental. For the 100M models, this peak is clearly visible around 8 layers. Models that are too shallow (e.g., 2-4 layers) or too deep (e.g., 12+ layers) for their parameter count are substantially less creative. Similarly, in Figure 3 (bottom), we visualize the impact of the width-to-depth ratio on the creativity of models at all three scales. Note that when depth is increased within a fixed parameter budget, the model’s width (embedding dimension) must necessarily decrease. For a fixed parameter count, there is also an optimal width-to-depth ratio that maximizes creativity, after which increasing the width further can be detrimental. The optimal $E/L$ ratio occurs between 200 and 300 for all three model sizes. This suggests that combinatorial creativity requires a delicate balance between (1) models that are too shallow and wide, where insufficient depth may hinder the sequential processing capacity to handle in-memory leaps of thought (which are required to make distant, constrained associations between concepts) and (2) models that are too deep and narrow, which suffer from restricted representational capacity that may limit their ability to hold and associate the diverse concepts needed for novel combinations.
@@
Realistic Asssumpt. Unrealistic assumptions
@@ 6 Limitations and Conclusion
While our work offers a promising theoretical framework for studying creativity, and our results offer exciting insights into the architectural choices that affect creativity, several limitations remain. Notably, we restricted our focus only to combinatorial creativity (CC), neglecting Boden (2004)’s other two forms (see Appendix C for additional commentary on this). Next, our empirical results relied on synthetic data, which may not be fully representative of the complexity of real-world data encountered in creative domains. Lastly, due to limited compute, we were only able to study up to 100M parameter models, whereas modern foundation models are well into the billions of parameters. Nevertheless, the generality of our framework means it is flexible enough to apply to real-world data, and future studies with access to more compute can explore the scaling behavior beyond the 100M cliff. Together, our conceptual framework and empirical findings offer a new pathway for understanding and improving the creativity of modern AI models, as we enter a new frontier in generalization abilities.
@@ C.1 Avenues for Improving Model Creativity
Appendix D LLM Statement
Large language models (LLMs) were used to generate the code used in this study, as well as for enhancing the clarity and flow of the text.

### v2

@@ Abstract
Artificial intelligence (AI) systems, and Large Language Models (LLMs) in particular, are increasingly employed for creative tasks like scientific idea generation, constituting a form of generalization from training data unaddressed by existing conceptual frameworks. Despite its similarities to compositional generalization (CG), combinatorial creativity (CC) is an open-ended ability. Instead of evaluating for accuracy or correctness against fixed targets, which would contradict the open-ended nature of CC, we propose a theoretical framework and algorithmic task for evaluating outputs by their degrees of novelty and utility. From here, we make several important empirical contributions: (1) We obtain the first insights into the scaling behavior of creativity for LLMs. (2) We discover that, for fixed compute budgets, there exist optimal model depths and widths for creative ability. (3) We find that the ideation-execution gap, whereby LLMs excel at generating novel scientific ideas but struggle to ensure their practical feasibility, may be explained by a more fundamental novelty-utility tradeoff characteristic of creativity algorithms in general. Importantly, this tradeoff remains persistent even at scale, casting doubt on the long-term creative potential of LLMs in their current form. Together, our conceptual framework and empirical findings provide a foundation for understanding and improving creativity in modern AI models, bridging the gap between human and machine intelligence.
[…]
@@ 1 Introduction
We now attempt to employ AI systems in scientifically creative tasks once conceptualized by Einstein (Gu and Krenn, 2024, Si et al., 2024, Sanyal et al., 2025), yet they lack strong mathematical and conceptual foundations for the abilities underlying these tasks. As a result, many problems have surfaced. LLM-generated ideas for scientific discovery often suffer from practical infeasibility, make unrealistic assumptions, and omit proper baselines, leading to what has been termed the ideation-execution gap (Si et al., 2025). Without a foundational understanding of creativity, our ability to diagnose and improve the outcomes of LLMs for such tasks remains severely limited.
To address these limitations, we introduce a formal framework and an open-ended, algorithmic task for evaluating combinatorial creativity. Our framework models creativity within a conceptual space, represented as a large graph where models must find novel paths between concepts while adhering to logical constraints. Using this testbed, we conduct a large-scale empirical study of decoder-only Transformers, which have become the de facto architecture for LLMs, by systematically varying their size, depth, and width to yield several key insights.
First, we obtain the first insights into the scaling behavior for creativity, demonstrating a predictable improvement in performance with increased model size and training compute. Second, we uncover a crucial architectural insight: for a fixed computational budget, there is an optimal model depth and width that maximizes creativity, beyond which performance diminishes. Third, we perform a detailed error analysis, which reveals that as task complexity increases, models are more likely to fail by violating utility constraints rather than by producing non-novel outputs. This is explained by our fourth and most critical finding: a fundamental novelty-utility tradeoff, which persists even as model scale increases. This suggests the ideation-execution gap may be an intrinsic challenge for current architectures—and not merely a problem of scale. Together, our conceptual framework and empirical findings offer a launchpad for understanding and improving the creativity of modern AI models, as we enter a new frontier in generalization abilities.
@@ 2.2 Distinguishing Combinatorial Creativity from Classical Forms of Generalization
[…]
@@ 3.2 Quantifying Degrees of Novelty
[…]
@@ 3.3 Quantifying Degrees of Utility
The utility function consists of three main parts: the terms $\left(1+\alpha_{I}|\mathcal{I}|\right)$ and $\left(1+\alpha_{X}|\mathcal{X}|\right)$ scale the utility function in proportion to the number of inclusion and exclusion constraints, respectively, while the indicator term ensures that artifacts obey these constraints and start and end at the correct nodes. The number of total constraints also allows us to informally define “utility levels:” level $1$ refers to unconstrained query ($I=\emptyset,X=\emptyset$), while levels $l>1$ each consist of l1l-1 constraints. In Section 5, we use these utility levels to demonstrate the novelty-utility tradeoff in creative LLM outputs. We also explain how the constraints can serve as minimal abstractions of failure modes encountered when using LLMs for scientifically creative idea generation.
@@ 3.4 Measuring Creativity
With this framework in place, we now investigate how creativity varies with architectural design and compute scaling.
@@ 4 Experiments
(a) Impact of depth on creativity.
(b) Impact of width on creativity.
Figure 3: The impact of width and depth on creativity. These heatmaps visualize the combinatorial creativity of models across three distinct parameter budgets (1M, 10M, and 100M). For each budget, the vertical axis represents the amount of training compute in FLOPs. The color intensity corresponds to the model’s creativity score, while the horizontal axis represents the number of layers $L$ (Figure 3(a)) or the width to depth ratio $E/L$ (Figure 3(b)). The contours reveal a clear, non-monotonic trend: in Figure 3(a), creativity improves as layers are added up to a certain point, after which performance declines, and in Figure 3(b), creativity improves as the width is increased up to a certain point, after which performance also declines. The optimal depth becomes more pronounced at larger scales, with the 100M models achieving peak creativity around 8 layers, while the optimal performance for width is at an $E/L$ ratio between 200 and 300.
@@
We are interested in how fundamental architectural choices influence the creativity of LLMs on the task defined in Section 3. For example, Nagarajan et al. (2025) recently found creative gains from changing the pre-training objective from next-token to multi-token prediction. In this study, we are especially curious whether creative limits present in smaller models also persist at scale, as this can provide insight into whether continued scaling of LLMs can yield even greater gains in creativity, or if architectural innovations are needed to support more open-ended behaviors (Hughes et al., 2024).
@@ 5 Results and Discussion
In Figure 3(a), we visualize the impact of the number of layers $L$ on combinatorial creativity across all three model sizes. Our most significant finding is that for a fixed parameter count, there is an architectural “sweet spot,” an optimal number of layers that maximizes creativity, after which increasing depth further can be detrimental. For the 100M models, this peak is clearly visible around 8 layers. Models that are too shallow (e.g., 2-4 layers) or too deep (e.g., 12+ layers) for their parameter count are substantially less creative. Similarly, in Figure 3(b), we visualize the impact of the width-to-depth ratio on the creativity of models at all three scales. Note that when depth is increased within a fixed parameter budget, the model’s width (embedding dimension) must necessarily decrease. For a fixed parameter count, there is also an optimal width-to-depth ratio that maximizes creativity, after which increasing the width further can be detrimental. The optimal $E/L$ ratio occurs between 200 and 300 for all three model sizes. This suggests that combinatorial creativity requires a delicate balance between (1) models that are too shallow and wide, where insufficient depth may hinder the sequential processing capacity to handle in-memory leaps of thought (which are required to make distant, constrained associations between concepts) and (2) models that are too deep and narrow, which suffer from restricted representational capacity that may limit their ability to hold and associate the diverse concepts needed for novel combinations.
@@
Realistic Assumpt. Unrealistic assumptions
@@ 6 Limitations and Conclusion
While our work offers a promising theoretical framework for studying creativity, and our results offer exciting insights into the architectural choices that affect creativity, several limitations remain. Notably, we restricted our focus only to combinatorial creativity (CC), neglecting Boden (2004)’s other two forms (see Appendix C for additional commentary on this). Next, our empirical results relied on synthetic data, which may not be fully representative of the complexity of real-world data encountered in creative domains. Lastly, due to limited compute, we were only able to study up to 100M parameter models, whereas modern foundation models are well into the billions of parameters. Nevertheless, the generality of our framework means it is flexible enough to apply to real-world data, and future studies with access to more compute can explore the scaling behavior beyond the 100M cliff. Together, our conceptual framework and empirical findings offer a new pathway for understanding and improving the creativity of modern AI models, bridging the gap between human and machine intelligence.